# Motion of a charged test particle around a static black hole in a monopole magnetic field

### v1

@@ Abstract
We study the motion of a charged test particle in the spacetime with a spherically symmetric black hole which is immersed in a monopole magnetic field. We show that the radial motion of the charged test particle is governed by completely the same equation as that in the case of no magnetic field. This result implies that the black hole will acquire the electric charge if it is surrounded by the collisionless plasma composed of protons and electrons which obey the Maxwell velocity distribution. The drastically different situation appears in the tangential motions of charged test particles due to the magnetic field. The trajectory of a charged test particle in the black hole with the magnetic field of the order of 10 Gauss near the black hole, is confined on a very thin cone as long as the specific angular momentum of the particle is not much larger than the gravitational radius of the black hole times the speed of light. This result leads to a possibility that a plasma lump can hover over the black hole and is very hot, in the monopole magnetic field.
preprint: OCU-PHYS-618preprint: AP-GR-206
@@ I Introduction
Active Galactic Nuclei (AGN) are considered to be deeply related to black holes. The AGN are phenomena in each of which a narrow region at the center of a galaxy emits strong electromagnetic radiation that is almost equal to the entire galactic emission. The source of the radiative energy is thought to be the gravitational energy released by mass accretion onto the supermassive black hole Brandford-etal ; Lynden-Bell ; Bardeen ; Rees . Very long relativistic jets of plasma from AGNs are observed and their energy sources are regarded as one of important issues of the black hole astrophysics. Long jets implies that they should be steady for long time which is equal to or longer than the length of the jet divided by the speed of light. The Brandford-Znajek processBZ-process in which the rotational energy of the black hole is released through the neighboring magnetic field is regarded as a promising candidate of the energy source of the relativistic jet.
@@
By those reasons mentioned above, in order to understand the physical situation around a black hole, it is crucially important to understand the motion of charged particles around the black hole immersed in electromagnetic field. In this paper, we study the motion of a charged test particle around a static black hole immersed in a monopole magnetic field. By using results obtained through this study, we investigate a possibility of the electrification of the black hole, and also make a comment on the temperature of the collisionless plasma surrounding the black hole. Here we should mention that Khan and Chen studied a charged test particle in the spacetime with a rotating black hole which is immersed in a split magnetic monopole field Khan-Chen . However, their interest is on the magnetic Penrose process, and very different from our present purposes.
@@ II Static black hole in a monopole electromagnetic field
In this paper, we focus on situations similar to the neighborhood of, for example, SgrA*, or the center of M87. The mass of SgrA* is equal to about $4\times 10^{6}M_{\odot}$ and the magnetic field in its neighborhood is estimated to be about 1010G EHT-MW-5 , whereas the mass of the black hole candidate in M87 is equal to about $6.5\times 10^{9}M_{\odot}$ and the magnetic field in its neighborhood is estimated be also about 1010G EHT-M87-5 . The magnetic charge $Q_{\rm M}$ which generates the magnetic field $B^{a}$ of the value 1010G at $r=Nr_{\rm g}$ is estimated as follows, where $N$ is a free parameter larger than one, and
@@
is the gravitational radius of the black hole. We have
@@ III Motion of a charged test particle in monopole field
where $\nabla_{b}$ and is the covariant derivative, whereas $F_{ab}$ is the electromagnetic field strength tensor defined as
@@ III.1 Conserved quantities
$u^{b}\nabla_{b}\Bigl\{\bigl[mu_{a}+q\left(A_{a}+\nabla_{a}\chi\right)\bigr]\xi^{a}\Bigr\}=q\left[{\cal L}_{\xi}\left(A_{a}+\nabla_{a}\chi\right)\right]u^{a}.$ (16)
If ${\cal L}_{\xi}\left(A_{b}+\nabla_{b}\chi\right)=0$ holds, this equation implies that $\bigl[mu_{a}+q\left(A_{a}+\nabla_{a}\chi\right)\bigr]\xi^{a}$ is conserved.
In the spacetime whose metric is given by Eq. (1), the time coordinate basis vector $(\partial/\partial t)^{a}$ is a Killing vector, and there is the $SO(3)$ isometry whose generators are three Killing vectors given below. By defining the Cartesian coordinates $(x,y,z)$ for three dimensional space as
@@
$L_{(x)}^{a}\nabla_{a}\chi_{(x)}$ $=-\frac{Q_{\rm M}}{\sin\theta}(1-\cos\theta)\cos\varphi,$ (21)
$L_{(y)}^{a}\nabla_{a}\chi_{(y)}$ $=-\frac{Q_{\rm M}}{\sin\theta}(1-\cos\theta)\sin\varphi$ (22)
@@ III.2 Motion in radial direction
Equation (39) is identical to the corresponding equation for the charged test particle in the Schwarzschild spacetime with no magnetic charge $Q_{\rm M}=0$ but with non-vanishing electric charge $Q_{\rm E}\neq 0$. No effect of the monopole magnetic field appears in the radial motion of the charged test particle, and hence, for example, the radius of ISCO is unchanged from that with no magnetic field. [^2] It should be noted that ISCO radius the rotating black hole immersed in the asymptotically homogeneous magnetic field can be very different from the case without magnetic field IHK2012 .
@@
must hold, we find that the allowed domain for the motion of the charged test particle is determined by the inequality
@@ III.3 Motion in angular directions
and hence $u_{\theta}=r^{2}\dfrac{d\theta}{d\tau}=0$ holds in the coordinate we adopt. The motion of the charged test particle is restricted in a hypersurface of constant polar angle
@@
where $e$ and $m_{\rm p}$ are the elementary charge and the mass of a proton, respectively. In the neighborhood of SgrA*, ${\cal Q}_{\rm M}$ of both the proton and the electron is very large. Thus, we find, from Eq. (52), that, in the neighborhood of SgrA* and in the neighborhood of the center in M87, $|\vartheta|\ll 1$ holds for the charged test particle of ${\cal K}^{2}$ which is much less than ${\cal Q}_{\rm M}^{2}$.
@@ IV Electric charge acquired by the black hole in monopole magnetic field
We assume charged test particles compose plasma which is in local thermal equilibrium, and the velocity of a charged test particle is given by the Maxwell distribution
@@
$e_{(t)\mu}$ $=\left(-\sqrt{f(r)},0,0,0\right),$ (54)
@@
We find from this equation that, in the case of ${\cal K}^{2}<3$, there is no extremum of $V(R;{\cal K}^{2},{\cal Q}_{\rm E}=0)$, and hence the test particles with ingoing initial velocity necessarily enters the black hole.
@@
Since the L.H.S. of this inequality is non-negative, the R.H.S. should be non-negative so that there exists ${\cal K}^{2}$ which satisfies this inequality. Since we focus on the case of $R\gg 1$, the following inequality should hold so that Eq. (67) holds.
@@
$c_{3}$ $=-\frac{1}{R^{2}}\left(1-\frac{1}{R}\right)\left[4-\frac{27}{R^{2}}\left(1-\frac{1}{R}\right)^{2}\right].$ (75)
By solving Eq. (71) with respect to ${\cal K}^{2}$, we obtain the following inequality
@@
By using Eq. (61), we rewriter $v_{(\theta)}$ and $v_{(\varphi)}$ as
@@
The test particle which initially moves outward will fall into the black hole only if it is bounded. As shown in Appendix A, the probability of a bounded particle initially moving outward for falling into the black hole is much less than that of a particle initially moving inward. Hence hereafter we consider only the probability $P$ of a particle initially moving inward for falling into the black hole. It is given as
@@
V Collisionless plasma hovering on the black hole
@@
Now, let us consider a lump of collisionless plasma bounded by the gravity of the central black hole, which is assumed to be far from the black hole,
@@
The energy density of the plasma is assumed to be much less than that of the magnetic field and hence too small to affect the spacetime geometry. Thus the charged particles in the plasma can be treated as test particles. In such a situation, the energy equation for a test particle, which is derived from the normalization condition of the four velocity, is approximately given as
@@
From the condition Eq. (87), we obtain
@@
The velocities of protons and electrons on the same orbit are equal to each other since they are governed by the gravity of the black hole. Thus the average of the kinetic energy of protons is about 1833 times larger than that of an electron.
@@
Let us consider a lump of collisionless plasma composed of protons and electrons, whose size is assumed to be, for example, the order of the gravitational radius $r_{\rm g}$. Its radial extent, $R_{\rm max}-R_{\rm min}$, is of order one and the tangential extent is $2R_{\rm min}\Theta=1$. Here note that the half apex angle $\vartheta$ of the cone on which the motion of the charged test particle is restricted is much less $\Theta$ in the situation of our interest, $10\lesssim R_{\rm min}<R_{\rm max}\lesssim 100$ (see Eqs. (51) and (52)). This means that the Larmor radius of the charged test particle is much less than the size of the plasma lump, since the Larmor radius is less than $R_{\rm max}\sin\vartheta$. Hence the random motions of the charged test particles in the plasma lump are Larmor one if the plasma is really collisionless. In this case, the time average of the kinetic energy estimated in Eq. (100) might be regarded as the “temperature” of the plasma lump. Here it is very remarkable fact that the kinetic energy of a charged test particle is equal to that of the “Keplerian” motion, since the energy equation of the charged test particle is given by Eq. (83). Hence, the temperature of the plasma lump can be very high in the neighborhood of the black hole in the monopole magnetic field. The origin of this temperature is the gravity.
@@
As mentioned, in the case that the plasma is composed of protons and electrons, their kinetic energies should be very different from each other due to the mass difference. Even if the velocity distribution of the protons and that of the electrons becomes Maxwell one with the same temperature by the thermalization through some plasma instability, the domain allowed for the motions of protons and those of electrons will be very different. Hence, the local thermal equilibrium seems to be unachievable in the plasma lump hovering on the black hole.
@@
It should be noted that the average velocity of electrons will be equal to that of protons in a plasma lump. This situation is very different from the assumed situation in the discussion of the electrification of a static black hole, where we have assumed that the velocity distribution of electrons and that of protons are Maxwell one with not so different temperatures from each other. Hence, we should note that lumps of collisionless plasma will not be suppliers of electric charge to the black hole.
@@
Here note that the global monopole magnetic field is not a necessary condition for the hovering motion. If the configuration of the magnetic field only in the domain of the motion of the charged test particle is locally similar to the monopole magnetic field, the charged test particle will hover on the black hole. In the stationary axi-symmetric magnetosphere, the neighborhood of the rotational symmetry axis of a rotating black hole may be the case (see for example, Fig. 3 in Ref. MG2004 ). Thus, in the realistic situation like as the neighborhood of SgrA*, the heating of a plasma lump may occur in the vicinity of the symmetry axis of rotating black hole.
@@ V.2 Collisionless condition
The two-body relaxation time for the plasma composed of protons and electrons is estimated as the time scale for which the accumulated changes in velocity through multiple two-body Coulomb scatterings becomes the same order of the initial velocity. The shortest time scale is given by the electron-electron scattering.
@@
By assuming the isotropy of the velocity distribution, we have
@@
We find from Fig. 1 that $R_{\rm min}/2<{\cal K}^{2}<R_{\rm max}/2$ holds if we focus on particles bounded within $R_{\rm min}<R<R_{\rm max}$. We are interested in the case of $10<R_{\rm min}<R_{\rm max}<100$, and hence we should assume $5<{\cal K}^{2}<50$. Then, from Eq. (52), we have ${\cal Q}_{\rm M}^{2}\gg{\cal K}^{2}$, and hence we may approximate $\sin\vartheta$ as
@@
$\frac{T_{\rm relax}}{T_{\rm Larmor}}=\frac{9\varepsilon_{0}^{2}\left(m_{\rm e}c^{2}\right)^{2}}{2\sqrt{3}\thinspace e^{4}}\times\frac{{\cal K}^{3}|{\cal Q}_{\rm M}|}{r_{\rm g}nR^{5}\ln(\lambda_{\rm D}/\ell)}$ (107)
@@
$=\frac{9\varepsilon_{0}^{2}c^{3}m_{\rm e}{\cal K}^{3}}{\sqrt{3}e^{3}nR^{5}\ln(\lambda_{\rm D}/\ell)}\left(\frac{N^{5}}{N-1}\right)^{1\over 2}\left(\frac{|q|}{e}\right)\left(\frac{m_{\rm e}}{m}\right)|B_{N}|$ (108)
@@
$n$ $\ll\frac{9\varepsilon_{0}^{2}c^{3}m_{\rm e}{\cal K}^{3}}{\sqrt{3}e^{3}R^{5}\ln(\lambda_{\rm D}/\ell)}\left(\frac{N^{5}}{N-1}\right)^{1\over 2}\left(\frac{|q|}{e}\right)\left(\frac{m_{\rm e}}{m}\right)|B_{N}|$
$\simeq\frac{3.04\times 10^{21}}{\ln(\lambda_{\rm D}/\ell)}\left(\frac{100r_{\rm g}}{r}\right)^{2}\left(\frac{10}{N}\right)^{2}\left(\frac{{\cal K}}{5}\right)^{3}\left(\frac{|q|}{e}\right)\left(\frac{m_{\rm e}}{m}\right)\left(\frac{|B_{N}|}{10^{-3}{\rm T}}\right)[{\rm m}^{-3}]$ (109)
holds, the plasma can be regarded to be collisionless, where ,in the last equality, we have assumed, $N\gg 1$. This condition is satisfied in the situations of our interest.

### v2

@@ Abstract
We study the motion of a charged test particle in the spacetime with a spherically symmetric black hole which is immersed in a monopole magnetic field. We show that the radial motion of the charged test particle is governed by completely the same equation as that in the case of no magnetic field. This result implies that the black hole will acquire the electric charge if it is surrounded by the collisionless plasma composed of protons and electrons which obey the Maxwell velocity distribution. The drastically different situation from no magnetic field case appears in the angular motions of charged test particles due to the magnetic field. The trajectory of a charged test particle around the black hole in the magnetic field of the order of 10 Gauss near the event horizon is confined on a very thin cone as long as the specific angular momentum of the particle is not much larger than the gravitational radius of the black hole times the speed of light. This result leads to a possibility that a plasma lump can hover over the black hole and is very hot, in the monopole magnetic field.
preprint: OCU-PHYS-618preprint: AP-GR-206preprint: NITEP 268
@@ I Introduction
Active Galactic Nuclei (AGN) are considered to be deeply related to black holes. The AGN are phenomena in each of which a narrow region at the center of a galaxy emits strong electromagnetic radiation that is almost equal to the entire galactic emission. The source of the radiative energy is thought to be the gravitational energy released by mass accretion onto the supermassive black hole Brandford-etal ; Lynden-Bell ; Bardeen ; Rees ; Antonucci . Very long relativistic jets of plasma from AGNs are observed and their energy sources are regarded as one of important issues of the black hole astrophysics. Long jets implies that they should be steady for a long time which is equal to or longer than the length of the jet divided by the speed of light. The Brandford-Znajek processBZ-process in which the rotational energy of the black hole is released through the neighboring magnetic field is regarded as a promising candidate of the energy source of the relativistic jet.
@@
By those reasons mentioned above, in order to understand the physical situation around a black hole, it is crucially important to understand the motion of charged particles around the black hole immersed in electromagnetic field. In this paper, we study the motion of a charged test particle around a static black hole immersed in a monopole magnetic field. By using results obtained through this study, we investigate a possibility of the electrification of the black hole, and also make a comment on the temperature of the collisionless plasma surrounding the black hole. Here we should mention that Khan and Chen studied a charged test particle in the spacetime with a rotating black hole which is immersed in a split magnetic monopole field Khan-Chen . However, their interest is on the magnetic Penrose process, and different from our present purposes.
@@ II Static black hole in a monopole electromagnetic field
In this paper, we focus on situations similar to the neighborhood of, for example, Sgr A*, or the center of M87. The mass of Sgr A* is about $4\times 10^{6}M_{\odot}$ and the magnetic field in its neighborhood is estimated to be about 1010G EHT-MW-5 , whereas the mass of the black hole candidate in M87 is equal to about $6.5\times 10^{9}M_{\odot}$ and the magnetic field in its neighborhood is estimated be also about 1010G EHT-M87-5 . The magnetic charge $Q_{\rm M}$ which generates the magnetic field $B^{a}$ of the value 1010G at $r=Nr_{\rm g}$ is estimated as follows, where $N$ is a free parameter larger than one, and
@@
is the gravitational radius of the black hole. We have the following magnetic field for static observers:
@@ III Motion of a charged test particle in monopole field
where $\nabla_{b}$ is the covariant derivative, whereas $F_{ab}$ is the electromagnetic field strength tensor defined as
@@ III.1 Conserved quantities
$u^{b}\nabla_{b}\Bigl\{\bigl[mu_{a}+q\left(A_{a}+\nabla_{a}\chi\right)\bigr]\xi^{a}\Bigr\}=qu^{a}{\cal L}_{\xi}\left(A_{a}+\nabla_{a}\chi\right).$ (16)
If ${\cal L}_{\xi}\left(A_{a}+\nabla_{a}\chi\right)=0$ holds, this equation implies that $\bigl[mu_{a}+q\left(A_{a}+\nabla_{a}\chi\right)\bigr]\xi^{a}$ is conserved.
In the spacetime whose metric is given by Eq. (1), the time coordinate basis vector $(\partial/\partial t)^{a}$ is a Killing vector, and there is the SO(3)(3) isometry whose generators are three Killing vectors given below. By defining the Cartesian coordinates $(x,y,z)$ for three-dimensional space as
@@
$L_{(x)}^{a}\nabla_{a}\chi_{(x)}$ $=-\frac{Q_{\rm M}}{4\pi\varepsilon_{0}c}\frac{\cos\varphi}{\sin\theta},$ (21)
$L_{(y)}^{a}\nabla_{a}\chi_{(y)}$ $=-\frac{Q_{\rm M}}{4\pi\varepsilon_{0}c}\frac{\sin\varphi}{\sin\theta}.$ (22)
@@ III.2 Motion in radial direction
Equation (39) is identical to the corresponding equation for the charged test particle in the Schwarzschild spacetime with no magnetic charge, $Q_{\rm M}=0$, but with non-vanishing electric charge, $Q_{\rm E}\neq 0$. No effect of the monopole magnetic field appears in the radial motion of the charged test particle, and hence, for example, the radius of ISCO is unchanged from that with no magnetic field. [^2] It should be noted that ISCO radius the rotating black hole immersed in the asymptotically homogeneous magnetic field can be very different from the case without magnetic field IHK2012 .
@@
must hold by the future directed property of the four-velocity, we find that the allowed domain for the motion of the charged test particle is determined by the inequality
@@ III.3 Motion in angular directions
and hence $u_{\theta}=r^{2}\dfrac{d\theta}{d\tau}=0$ holds in the coordinates we adopt. The motion of the charged test particle is restricted in a hypersurface of constant polar angle
@@
where $e$ and $m_{\rm p}$ are the elementary charge and the mass of a proton, respectively. In the neighborhood of Sgr A*, ${\cal Q}_{\rm M}$ of both the proton and the electron is very large. Thus, we find, from Eq. (52), that, in the neighborhood of Sgr A* and in the neighborhood of the center in M87 galaxy, $|\vartheta|\ll 1$ holds for the charged test particle of ${\cal K}^{2}$ which is much less than ${\cal Q}_{\rm M}^{2}$.
@@ IV Electric charge acquired by the black hole in monopole magnetic field
We assume that charged test particles compose plasma which is in local thermal equilibrium, and that the velocity of a charged test particle is given by the Maxwell distribution
@@
$e_{(t)\mu}$ $=\left(-c\sqrt{f(r)},0,0,0\right),$ (54)
@@
We find from this equation that, in the case of ${\cal K}^{2}<3$, there is no extremum of $V(R;{\cal K}^{2},{\cal Q}_{\rm E}=0)$, and hence the test particles with ingoing initial velocity necessarily enter the black hole.
@@
Since the L.H.S. of this inequality is non-negative, the R.H.S. should be non-negative so that there exists ${\cal K}^{2}$ which satisfies this inequality. Since we focus on the case of $R\gg 1$ initially, the following inequality should hold so that Eq. (69) holds:
@@
$c_{3}$ $=-\frac{1}{R^{2}}\left(1-\frac{1}{R}\right)\left[4-\frac{27}{R^{2}}\left(1-\frac{1}{R}\right)\right].$ (75)
By solving Eq. (71) with respect to ${\cal K}^{2}$, we obtain the following inequality:
@@
By using Eq. (61), we rewrite $v_{(\theta)}$ and $v_{(\varphi)}$ as
@@
A test particle which initially moves outward will fall into the black hole only if it is bounded. As shown in Appendix A, the probability of a bounded particle initially moving outward for falling into the black hole is much less than that of a particle initially moving inward. Hence hereafter we consider only the probability $P$ of a particle initially moving inward for falling into the black hole. It is given as
@@
V Collisionless plasma hovering above the black hole
@@
Now, let us consider a lump of collisionless plasma bounded by gravity of the central black hole, which is assumed to be far from the black hole in the range,
@@
The energy density of the plasma is assumed to be much less than that of the magnetic field and hence too small to affect the spacetime geometry. Thus the charged particles in the plasma can be treated as test particles. In such a situation, the energy equation for a test particle, which is derived from the normalization condition of the four-velocity, is approximately given as
@@
From the condition Eq. (87), (88) and (90), we obtain
@@
The velocities of protons and electrons on the same orbit are equal to each other since they are governed by gravity of the black hole. Thus the average of the kinetic energy of protons is about 1833 times larger than that of an electron.
@@
Let us consider a lump of collisionless plasma composed of protons and electrons, whose size is assumed to be, for example, the order of the gravitational radius $r_{\rm g}$. Its radial extent, $R_{\rm max}-R_{\rm min}$, is of order one and the tangential extent is $2R_{\rm min}\Theta=1$. Here note that the half apex angle $\vartheta$ of the cone on which the motion of the charged test particle is restricted is much less than $\Theta$ in the situation of our interest, $10\lesssim R_{\rm min}<R_{\rm max}\lesssim 100$ (see Eqs. (51) and (52)). This means that the Larmor radius of the charged test particle is much less than the size of the plasma lump, since the Larmor radius is less than $R_{\rm max}\sin\vartheta$. Hence the random motions of the charged test particles in the plasma lump are Larmor one if the plasma is really collisionless. In this case, the time average of the kinetic energy estimated in Eq. (100) might be regarded as the “temperature” of the plasma lump. Here it is very remarkable fact that the kinetic energy of a charged test particle is equal to that of the “Keplerian” motion, since the energy equation of the charged test particle is given by Eq. (83). Hence, the temperature of the plasma lump can be very high in the neighborhood of the black hole in the monopole magnetic field. The origin of this temperature is gravity.
@@
As mentioned, in the case that the plasma is composed of protons and electrons, their kinetic energies should be very different from each other due to the mass difference. Even if the velocity distribution of the protons and that of the electrons become Maxwell ones with the same temperature, the domain allowed for the motions of protons and those of electrons will be very different. Hence, the local thermal equilibrium seems to be unachievable in the plasma lump hovering on the black hole.
@@
It should be noted that the average velocity of electrons will be equal to that of protons in a plasma lump. This situation is very different from the situation assumed in the discussion of the electrification of a static black hole of the NHYI paper paper-I , where we have assumed that the velocity distribution of electrons and that of protons are Maxwell one with not so different temperatures from each other. Hence, we should note that lumps of collisionless plasma will not be suppliers of electric charge to the black hole.
@@
Here note that the global monopole magnetic field is not a necessary condition for the hovering motion. If the configuration of the magnetic field only in the domain of the motion of the charged test particle is locally similar to the monopole magnetic field, the charged test particle will hover on the black hole. In the stationary axisymmetric magnetosphere, the neighborhood of the rotational symmetry axis of a rotating black hole may be the case (see for example, Fig. 3 in Ref. MG2004 ). Thus, in the realistic situation like the neighborhood of Sgr A*, the heating of a plasma lump may occur in the vicinity of the symmetry axis of rotating black hole.
@@ V.2 Collisionless condition
The two-body relaxation time for the plasma composed of protons and electrons is estimated as the time scale for which the accumulated changes in velocity through multiple two-body Coulomb scatterings becomes the same order of the initial velocity. The shortest time scale is given by the electron-electron scattering,
@@
By assuming the isotropy of the velocity distribution, we have, form Eq. (31),
@@
We find from Fig. 1 that $R_{\rm min}/2<{\cal K}^{2}<R_{\rm max}/2$ holds if we focus on particles bounded within $R_{\rm min}<R<R_{\rm max}$. We consider the case of $10<R_{\rm min}<R_{\rm max}<100$, and hence we should assume $5<{\cal K}^{2}<50$. Then, from Eq. (52), we have ${\cal Q}_{\rm M}^{2}\gg{\cal K}^{2}$, and hence we may approximate $\sin\vartheta$ as
@@
$=\frac{9\varepsilon_{0}^{2}c^{3}m_{\rm e}{\cal K}^{3}}{2\sqrt{3}e^{3}nR^{5}\ln(\lambda_{\rm D}/\ell)}\left(\frac{N^{5}}{N-1}\right)^{1\over 2}\left(\frac{|q|}{e}\right)\left(\frac{m_{\rm e}}{m}\right)|B_{N}|,$ (107)
where we have used Eq. (38) with Eq. (7) in the second equality.
@@
$n$ $\ll\frac{9\varepsilon_{0}^{2}c^{3}m_{\rm e}{\cal K}^{3}}{2\sqrt{3}e^{3}R^{5}\ln(\lambda_{\rm D}/\ell)}\left(\frac{N^{5}}{N-1}\right)^{1\over 2}\left(\frac{|q|}{e}\right)\left(\frac{m_{\rm e}}{m}\right)|B_{N}|$
$\simeq\frac{1.52\times 10^{22}}{\ln(\lambda_{\rm D}/\ell)}\left(\frac{100r_{\rm g}}{r}\right)^{5}\left(\frac{N}{10}\right)^{2}\left(\frac{{\cal K}}{5}\right)^{3}\left(\frac{|q|}{e}\right)\left(\frac{m_{\rm e}}{m}\right)\left(\frac{|B_{N}|}{10^{-3}{\rm T}}\right)[{\rm m}^{-3}]$ (108)
holds, the plasma can be regarded to be collisionless, where, in the last equality, we have assumed $N\gg 1$. This condition is satisfied in the situations of our interest.